the sum of the ages of a father and son is 46 years. In 2 years, four times the sun's age will equal the fathers age. What are their present ages?
Answer by macston(5194)

F=father's age; S=son's age
F+S=46 years
F=46 years-S
F+2=4(S+2) Substitute for F
(46 years-S)+2=4S+8 Add S to each side
48 years=5S+8 Subtract 8 from each side.
40 years=5S Divide each side by 5.
8 years=S ANSWER 1: The son is 8 years old.
F=46 years-S=46 years-8 years=38 years ANSWER 2: Father is 38 years old.
CHECK:
F+2=4(S+2)
38 years + 2 years=4(8 years+2 years)
40 years=4(10 years)
40 years=40 years